This is an actual question on a ACT practice test. their answer is 3.08 How did they arrive at this answer?
In a college, some courses contribute more towards an overall GPA than other courses. For example, a science class is worth 4 points; mathematics is worth 3 points; History is worth 2 points; and English is worth 3 points. The values of the grade letters are as follows, A= 4, B=3, C=2, D=1, F=0. What is the GPA of a student who made a “C” in Trigonometry, a “B” in American History, an “A” in Botany, and a “B” in Microbiology?

2.59
2.86
3.08
3.33
3.67

Answer by Earlsdon(6098)

GPA (Grade Point Average) is computed by:
GPA = Total # of grade points divided by Total number of course points.
Botany (Science = 4 course points) = A (4X4 grade points = 16)
Microbiology (Science = 4 course points) = B (4X3 grade points = 12)
Trigonometry (Math = 3 course points) = C (3X2 grade points = 6)
American History (History = 2 course points) = B (2X3 grade points = 6)
Total course points = 13.
Total grade points = 40.
GPA = 40/13
GPA = 3.0769
GPA = 3.08 Rounded to nearest hundredth.